The problem is, "The length of a rectangle exceeds it's width by 3 meters. The perimeter is 58 meters. Find the dimensions." I tried using subsitutions, but I got stuck and didn't get an answer. What do I do?
Answer by ankor@dixie-net.com(22740)

"The length of a rectangle exceeds it's width by 3 meters.
The perimeter is 58 meters. Find the dimensions."
:
The perimeter: 2L + 2W = 58
We can simplify this, divide by 2, and you have
L + W = 29
:
We know: L = (W+3)
:
(W+3) + W = 29
2W = 29 - 3
2W = 26
W = 13 is the width
then
13 + 3 = 16 is the Length
;
Now you check this in the perimeter equation